OpenStudy (anonymous):

Using complete sentences, explain how to graph y = one thirdx + 4 using the slope-intercept method.

OpenStudy (anonymous):

You can get in trouble for medal fishing

OpenStudy (ranga):

y = 1/3 * x + 4 Slope = 1/3 and y-intercept = 4. y-intercept is the point on the y-axis where the line crosses the y-axis. So mark the point (0,4) on the y-axis. That is one point. We just need one more point on the line to draw the line. The slope is 1/3. Slope = rise / run. Start from the point (0,4) we found earlier. Move 1 unit above along the y-axis and move 3 units horizontally. The point will be (0+3, 4+1) = (3,5). Mark that point. Using a ruler drawing a straight line through the points (0,4) and (3,5).